OpenStudy (anonymous):

what is a quadratic function in the form of y = ax2 + bx + c that would have the given roots? 1. x = 2i and x = –2i 2. x =3 + i and x = 3 – i

OpenStudy (inkyvoyd):

hint.

OpenStudy (inkyvoyd):

quaderatic equation with roots alpha and beta

OpenStudy (inkyvoyd):

can be written as (x-alpha)(x-beta)=0

OpenStudy (anonymous):

1- (x+2i)(x-2i) = x^2 -4i^2 = x^2 +4 2- (x- (3+i)) (x- (3-i)) = x^2 + (3+i)(3-i) = x^2 + (9 -i^2) = x^2 - 9

OpenStudy (anonymous):

1) (x-2i)(x+2i) Multiply it out: x^2 + 2ix - 2ix - 4i^2 Simplify: x^2+4 If you really need a, b, and c in there: y = 1x^2 + 0x + 4 2) [x-(3+i)][x-(3-i)] Multiply it out: x^2 - (3-i)x - (3+i)x + (3+i)(3-i) x^2 - 6x - ix + ix + 9 - 3i + 3i - i^2 Simplify: x^2 - 6x + 9 + 1 x^2 + 6x + 10 y = 1x^2 + 6x + 10 this is my answer but i think i did it wrong -___-

OpenStudy (anonymous):

The first one is right as I get the same answer but the second one I guess I did something wrong I think you're right in the second one

OpenStudy (anonymous):

@julieteatsyou

OpenStudy (anonymous):